OpenStudy (anonymous):

how do i put this in scientific notation ? (2.9 x 10^3)(3.0 x 10^-3)

OpenStudy (anonymous):

2.9 * 3 = 8.7 10^3 * 10^-3 = 10 ^0 = 1 8.7 x 10^0 = 8.7